if i add my current age to the age i will be by next year on this date, the sum is 103 years. how old will i be 10 years from today?
Answer by Kalmetam(43)

you have to be around the half of 102 area and half of 102 is 51. Lets try.
50+51=101 X
51+52=103 O
52+53=105 X
So now you have to be 51. 51 plus ten is obviously 61 so you will be 61 in ten years.